Calculus I, Fall 2015

Quiz 3

Nov 12, 2015 Show all details.

1. Find an approximate value of \sqrt[3]{1.009} and show that the error of the approximation is smaller than 10^{−5}.

2. Show that, if Rolle’s Theorem is true, then the Mean Value Theorem is true.

3. Let f(x) = x^{2/3}(3 − x). Find absolute max and min of f on [−3, 3]. Show all details.

4. Graph y = \frac{1}{x^{2}−1}. Show all details.

Calculus I, Fall 2015

Quiz 3

Nov 12, 2015 Show all details.

1. Find an approximate value of \sqrt[3]{1.009} and show that the error of the approximation is smaller than 10^{−5}.

2. Show that, if Rolle’s Theorem is true, then the Mean Value Theorem is true.

3. Let f(x) = x^{2/3}(3 − x). Find absolute max and min of f on [−3, 3]. Give all details.

4. Graph y = \frac{1}{x^{2}−1}. Give all details.